OpenStudy (anonymous):

3x-7y=21. Need the 3 ordered pairs where #1: x=0, is __ #2: y=0, is __ #3: y=-6, is __.

OpenStudy (anonymous):

ok, so for the first one, you would plug in 0( which is x) in the equation.: 3(0)-7y=21 0-7y=21 y=21/-7 y=-3, so first ordered pair would be (0, -7) for the second one you would plug in0 for y this time, and in the third one you would plug in -6 for why, and solve for x on that too. then you have your three ordered pairs

OpenStudy (anonymous):

I am still not getting it so need more help

OpenStudy (anonymous):

ok, so lets do the second one

OpenStudy (anonymous):

ok i tried but got lost

OpenStudy (anonymous):

the second one says y= 0 so we are going to plug in 0 for the y, in the function: 3x-7y=21 so when we plug in 0 for y, the function now says: 3x-7(0)=21 next we multiply 0 times 7, so we get: 3x-0=21, which is basically 3x=21 next, we finaly divide 3 on both sides to get x by itself x=21/3 which is x =7

OpenStudy (anonymous):

so now that we know that x=7 , when y=0, the ordered pair is in the form of (x,y) which in this case, it would be (7, 0). that is your second set of points. did you understand this?

OpenStudy (anonymous):

well when i did it it was a little different but it did not look right but had the same answer

OpenStudy (anonymous):

for the last one i started with 3x-7(-6)=21

OpenStudy (anonymous):

then i had 3x-42=21

OpenStudy (anonymous):

ok wait a sec

OpenStudy (anonymous):

whats -7 *-6?

OpenStudy (anonymous):

don't I replace the y with -6?

OpenStudy (anonymous):

yes, but what is -7 times -6. can you hold up for a sec

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

3x-7(-6)=21 is the correct next step, but what do you get when you multiply negative 7 times negative 6?

OpenStudy (anonymous):

you get positive 42

OpenStudy (anonymous):

so you should have 3x-42=21

OpenStudy (anonymous):

no, its going to be 3x+42=21 because we got positive 42, not negative 42

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

yes 3x+42=21, what do we do next?

OpenStudy (anonymous):

get x by itself by using 3

OpenStudy (anonymous):

OK WAIT

OpenStudy (anonymous):

we still have that 42 there, and we need to get rid of it, so how do we get rid of that 42: 3x+42=21

OpenStudy (anonymous):

3x +42 -42 = 21 +42 - 42

OpenStudy (anonymous):

there we go! and what do we get?

OpenStudy (anonymous):

3x=21

OpenStudy (anonymous):

wait whats 21-42? -21!

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

I thought the 42s cancelled each other out

OpenStudy (anonymous):

yes, but its not 3x=21 its 3x=-21

OpenStudy (anonymous):

ok I don't get it

OpenStudy (anonymous):

so this is what you got 3x +42 -42 = 21 +42 - 42

OpenStudy (anonymous):

right

OpenStudy (anonymous):

why did you add 42 both sides? you were only supposed to subtract 42 on both sides. try it again,

OpenStudy (anonymous):

3x +42 -42 = 21- 42 is what you should get

OpenStudy (anonymous):

ok i put an extra one in there

OpenStudy (anonymous):

exactly so figure it from this: 3x +42 -42 = 21- 42

OpenStudy (anonymous):

ok so 3x +42 -42 = 0

OpenStudy (anonymous):

21 -42= -21

OpenStudy (anonymous):

yes so it should be 3x=-21

OpenStudy (anonymous):

so (-21,-6) is the ordered pair

OpenStudy (anonymous):

we stil need to find x! we were left with 3x=-21 ! remember we need to get x by itself!

OpenStudy (anonymous):

ok so we us 3 on both sides

OpenStudy (anonymous):

we divide 3 by both sides

OpenStudy (anonymous):

damn this thing hangs up on me a lot. so we should end up with x = 21

OpenStudy (anonymous):

whats -21/3?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

so x=-7. gosh, what grade are you in? and math level?

OpenStudy (anonymous):

I just started college last year Oct and I graduated in 1969 and had just basic math. I was able to transfer 52 credits from my military career. So I am about a late sophmore or early junior. math is 092 which is pre algebra

OpenStudy (anonymous):

college 0.0i learned this in 6th and 7th.

OpenStudy (anonymous):

ya I know but I have never had algebra before

OpenStudy (anonymous):

and im in 9th. yaaay! i feel so smart!

OpenStudy (anonymous):

when i was in high school we did not have algebra and geometry. it did not exist

OpenStudy (anonymous):

so when i was in 12th they were just starting it as an elective.

OpenStudy (anonymous):

I guess that is why it is so hard for me.

OpenStudy (anonymous):

whoa! it really does amaze me how far human intelligence has gone. I cant imagine how advanced our human race will be in he next 40 years

OpenStudy (anonymous):

ya i know. i wish there was a way for me to shock the heck out my brain so it would be easier for me to learn

OpenStudy (anonymous):

lol, so do i! agh sooo much work!

OpenStudy (anonymous):

ya but you understand this. I son't understand any of it. I just started in Jan to learn this junk

OpenStudy (anonymous):

ya, this si easy. but im doing precal, and ap microecon, and ap Human Geogarohy, and ap bio